OpenStudy (anonymous):

All read numbers that are less than -1 or greater tha or equal to 4

OpenStudy (anonymous):

inequality or interval notation?

OpenStudy (anonymous):

you could write \[x<-1\text{ or } x\geq 4\]

OpenStudy (anonymous):

you could write \[(-\infty, -1)\cup [4,\infty)\]